# TLOB: A Novel Transformer Model with Dual Attention for Stock Price Trend Prediction with Limit Order Book Data

### v2

@@ Abstract.
Stock Price Trend Prediction (SPTP) based on Limit Order Book (LOB) data is a fundamental challenge in financial markets. Despite advances in deep learning, existing models fail to generalize across different market conditions and struggle to reliably predict short-term trends. Surprisingly, by adapting a simple MLP-based architecture to LOB, we show that we surpass SoTA performance; thus, challenging the necessity of complex architectures. Unlike past work that shows robustness issues, we propose TLOB, a transformer-based model that uses a dual attention mechanism to capture spatial and temporal dependencies in LOB data. This allows it to adaptively focus on the market microstructure, making it particularly effective for longer-horizon predictions and volatile market conditions. We also introduce a new labeling method that improves on previous ones, removing the horizon bias. We evaluate TLOB’s effectiveness using the established FI-2010 benchmark, which exceeds the state-of-the-art by an average of 3.7 F1-score(%). Additionally, TLOB shows improvements on Tesla and Intel with a 1.3 and 7.7 increase in F1-score(%), respectively. Additionally, we empirically show how stock price predictability has declined over time (-6.68 absolute points in F1-score(%)), highlighting the growing market efficiencies. Predictability must be considered in relation to transaction costs, so we experimented with defining trends using an average spread, reflecting the primary transaction cost. The resulting performance deterioration underscores the complexity of translating trend classification into profitable trading strategies. We argue that our work provides new insights into the evolving landscape of stock price trend prediction and sets a strong foundation for future advancements in financial AI. We release the code at https://github.com/LeonardoBerti00/TLOB.
Stock Price Prediction, Limit Order Book, Transformer
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; August 03–07, 2025; Toronto, Canadaccs: Computing methodologies Neural networks
1. Introduction
@@
Over the past few decades, the global financial landscape has undergone a profound transformation, transitioning from manual trading operations to sophisticated electronic platforms. This evolution has been so significant that by 2020, electronic trading accounted for over 99% of equity shares traded in the United States, a stark contrast to just 15% in 2000 (Kissell, 2020). At the heart of this revolution lies the electronic Limit Order Book (LOB), a dynamic data structure that has become the cornerstone of modern financial markets. In today’s competitive financial world, the majority of the markets utilize electronic LOBs to record trades. The continuous inflow of limit orders, organized by price levels, creates a dynamic structure that evolves over time, reflecting the real-time balance of supply and demand. However, this multidimensional structure, which spans price levels and volumes, presents complex challenges for understanding market behavior, forecasting stock price trends, and simulating realistic market conditions. The non-stationary nature of LOB data, characterized by its stochastic behavior, makes modeling stock price movements challenging. Traditional statistical methods fail to capture these complexities, especially when attempting to predict short-term price trends. However, recent advancements in deep learning have opened new avenues for tackling these challenges, offering the ability to model the non-linear relationships and temporal dependencies inherent in LOB data.
@@
Stock Price Trend Prediction (SPTP) [^1] In the literature it is also referred to as mid-price movement prediction. remains one of the most challenging and economically significant problems in financial markets, attracting significant attention from academic researchers and industry practitioners. One prominent application of SPTP, particularly utilizing Limit Order Book (LOB) data, lies within high-frequency trading, where algorithms attempt to capitalize on short-term price movements. Predicting future market movements is a highly challenging task due to the complexity, non-stationarity, and volatility of financial markets. However, with the growing availability of Limit Order Book (LOB) data and advancements in deep learning, new opportunities have emerged to improve the accuracy of these predictions. This paper explores the application of deep learning models to SPTP using Limit Order Book (LOB) data, which provides the most granular and complete information on stock trades. Financial markets do not exist in a vacuum; they are continuously shaped by the actions and expectations of countless participants who, according to the Efficient Market Hypothesis (EMH), collectively incorporate all available information into asset prices. When models discover a predictive pattern and traders act on it, the anomaly is quickly competed away, causing a paradox: successful signals sow the seeds of their own demise. Over time, greater liquidity, advanced trading technologies, and the proliferation of algorithmic strategies intensify this effect, i.e., any exploitable signal becomes visible in execution data and erodes more rapidly. Consequently, the apparent decline in forecast accuracy in our findings aligns with EMH principles: as soon as new patterns are detected and exploited, the relentless engine of arbitrage drives markets back toward efficiency. This interplay underlines why forecasting often becomes more difficult the farther we move from idealized, less liquid markets (like FI-2010) toward active, high-efficiency markets (like NASDAQ), thereby illustrating a core tension between the pursuit of alpha and the self-correcting nature of competitive markets. Traditional approaches relied on technical analysis and statistical methods, but recent years have seen a shift toward more sophisticated deep learning methods. A lot of different types of deep learning architectures have been utilized to tackle the SPTP tasks. Tsantekidis et al. utilized Long Short-Term Memory (LSTM) layers (Tsantekidis et al., 2017a) and Convolutional Neural Networks (CNNs) (Tsantekidis et al., 2017b, 2020). Zhang et al. (Zhang et al., 2019) introduced the DeepLOB model, which leverages LOB data to predict mid-price movements using a combination of convolutional and LSTM layers. Recent work (Prata et al., 2024) has highlighted the limitations of existing models, particularly their lack of robustness and generalizability when applied to diverse market conditions and more efficient stocks. In this paper, we address these limitations by proposing TLOB, a transformer-based approach, that outperforms all the existing models on both benchmark and real-world datasets, paving the way for more reliable SPTP applications. We introduce also an MLP-based model to show that a simple architecture, based on fully connected layers and GeLU activation function, can outperform all the SoTA models. We list our contributions:
  1. (1)
@@
    Novel Architecture Proposals: We introduce two new deep learning models:
@@
  2. (2)
    Comprehensive Evaluation: We conduct extensive experiments on both the benchmark FI-2010 dataset and a real-world NASDAQ dataset composed of Tesla and Intel stocks, with several baselines, providing insights into model performance across different market conditions and time horizons. We also perform an ablation study investigating the design choices of TLOB.
  3. (3)
@@ 2. Background
The LOB is a data structure that maintains and matches active limit orders and market orders in accordance with a predefined set of rules. This structure is transparently accessible to all market participants and is subject to continuous updates with each event, including order placement, modification, cancellation, and execution. The most widely adopted mechanism for order matching is the Continuous Double Auction (CDA) (Bouchaud et al., 2018). Under the CDA framework, orders are executed whenever the best bid (the highest price a buyer is willing to offer) and the best ask (the lowest price a seller is willing to accept) overlap. This mechanism facilitates continuous and competitive trading among market participants. The price of a security is commonly defined as the mid-price, calculated as the average of the best ask and best bid prices, with the difference between these prices representing the bid-ask spread.
@@
Given that limit orders are organized into distinct depth levels, each comprising bid price, bid size, ask price, and ask size, based on their respective prices, the temporal evolution of a LOB constitutes a complex, multidimensional temporal problem. Research on LOB data can be broadly categorized into four primary types: empirical analyses of LOB dynamics (Cont, 2001; Bouchaud et al., 2002), price and volatility forecasting (Zhang et al., 2019; Sirignano, 2019), stochastic modeling of LOB dynamics (Cont, 2011; Gould et al., 2013), and LOB market simulation (Byrd et al., 2020; Coletta et al., 2021; Li et al., 2020).
3. Related Work
The challenge of modeling the complex data structures and vast quantities associated with LOBs has spurred the development of deep learning algorithms for related modeling and forecasting tasks. In this section, we will summarize the State-of-The-Art (SoTA) deep learning models in the Stock Price Trend Prediction (SPTP) task, which consists of forecasting the direction of mid-price movements at a high-frequency resolution. Tsantekis et al.(Tsantekidis et al., 2017b) (2017) utilize a Recurrent Neural Network (RNN) based on Long-Short Term Memory (LSTM) layers to predict mid-price movements. In the same year, the authors presented another approach (Tsantekidis et al., 2017a), introducing a CNN-based model (CNN). Subsequently, the same group proposed two additional architectures in (Tsantekidis et al., 2020) (2020). The first focuses on capturing temporal dynamics from LOB data and correlating temporally distant features using convolutional layers. The second architecture, CNNLSTM, merges the CNN with an LSTM. The CNN initially extracts features from the LOB time series, which are then passed to the LSTM for classification. Tran et al.(Tran et al., 2018) (2018) proposed the Temporal Attention-Augmented Bilinear Layer (TABL) for multivariate time series prediction. This architecture, applies a bilinear transformation to the input, capturing dependencies between features and over time. The authors extended this model in (Tran et al., 2021), introducing BINCTABL, which integrates a bilinear normalization layer into the architecture to address non-stationarity and magnitude disparity (for instance prices and sizes) within the time series data. Passalis et al.(Passalis et al., 2019) (2019) introduced DAIN (Deep Adaptive Input Normalization), a three-step layer that adaptively normalizes data, instead of relying on fixed precomputed statistics. DAIN comprises three layers: a shifting layer, a scaling layer, and a gating layer, which suppresses irrelevant features by applying a sigmoid function. DAIN was integrated into various architectures, including MLPs (Nousi et al., 2019), CNNs (Tsantekidis et al., 2017a), and RNNs (Cho, 2014). Zhang et al. (Zhang et al., 2019) (2019) introduced DEEPLOB, which consists of three main blocks: convolutional layers, an Inception Module, and an LSTM layer. The convolutional layers and Inception Module extract relevant features, while the LSTM captures temporal dependencies. Two years later, Zhang et al.(Zhang and Zohren, 2021) (2021) extended DEEPLOB by adopting the attention (Luong, 2015) mechanism, creating DEEPLOBATT for multi-horizon forecasting. In this architecture, an encoder extracts features from LOB data, and an attention mechanism assigns weights to the hidden states, improving the processing of long input sequences. Finally, Kiesel et al.(Kisiel and Gorse, 2022) (2022) introduced Axial-LOB, which uses axial attention (Ho et al., 2019) to factorize 2D attention into two 1D attention modules, one for the feature axis and one for the time axis. Prata et al. (Prata et al., 2024) evaluated 15 deep learning models for stock prediction using limit order book data, including all the models described above. Some performed well on the FI-2010 dataset, but most of them showed non-reproducible results. When trained and tested on a new dataset composed of NASDAQ stocks most failed to generalize, in particular the performances of every model flattened around 60 in F1-Score. BiN-CTABL ((Tran et al., 2021)) was the top performer, and attention-based models generally excelled. The main reason given by the authors is that NASDAQ stocks are more complex to forecast than Finnish ones and the SoTA models cannot capture this complexity. Furthermore, all models showed high sensitivity to hyperparameters and context. These results make the models unreliable for real-world use.
Another stream of research has focused on meta-learning Transformer models, i.e., TabPFN models (Hollmann et al., 2022; Hoo et al., 2025), that excel at “one forward-pass” inference on small tabular datasets, leveraging massive synthetic corpora of low-dimensional tasks to learn prior feature distributions. However, they become computationally prohibitive for large-scale LOB data, where millions of high-dimensional observations strain both memory and processing. Consequently, TabPFN-based methods remain impractical for demanding real-world predictive tasks, such as SPTP, underscoring the need for more scalable deep learning architectures (e.g., LSTM, CNN, or specialized Transformers) tailored to extensive time series.
Figure 1. Comparison of three labeling methods. $t$ is the current timestamp, $k$ is the smoothing window length, and $h$ is the prediction horizon. In our proposed method (c), $k$ and $h$ are defined independently, providing a more flexible and unbiased approach.
4. Task Definition
Figure 2. TLOB architecture overview. The model leverages Temporal Self-Attention and Feature Self-Attention within each TLOB block to capture time-wise and spatial relationships in Limit Order Book data. Each block is preceded by Bilinear Normalization to address non-stationarity, followed by an MLPLOB block.
@@ 4. Task Definition
(1) $\mathbb{L}(t)=\bigl(P^{ask}(t),\,V^{ask}(t),\,P^{bid}(t),\,V^{bid}(t)\bigr),$
where $P^{ask}(t)$ and $P^{bid}(t)\in\mathbb{R}^{L}$ are the prices at levels $1$ through $L$, and $V^{ask}(t)$ and $V^{bid}(t)\in\mathbb{R}^{L}$ are the corresponding volumes. Trend Definition We employ a ternary classification system for price trends: U (“upward”) denotes an increasing price trend, D (“downward”) indicates a decreasing trend, and S (“stable”) represents price movements with only minor variations.
In equity markets, mid-prices are generally considered the most reliable single-value indicator of actual stock prices. However, owing to inherent market fluctuations and exogenous shocks, mid-prices can exhibit considerable volatility. Consequently, labeling consecutive mid-prices $\bigl(p_{t},p_{t+1}\bigr)$ often results in noisy labels.
To mitigate this, many labeling strategies employ smoother mid-price functions, averaging prices over a chosen “window length” to reduce short-term noise and better reflect persistent directional moves. An example of this approach appears in (Ntakaris et al., [n. d.]), detailed in Section 6.2.
However, as shown by Zhang et al. (Zhang et al., 2019) (Fig. 2), smoothing only the future prices can lead to instability in trading signals. This instability often causes redundant trading actions and higher transaction costs. To address this, Tsantekidis et al. (Tsantekidis et al., 2017a) proposed also smoothing past prices. They define:
(2) $l(t,k)\,=\,\frac{m_{+}(t,k)\;-\;m_{-}(t,k)}{m_{-}(t,k)}\quad\text{where}$
(3) $m_{+}\bigl(t,k\bigr)\;=\;\frac{1}{k+1}\sum_{i=0}^{k}p\bigl(t+i\bigr)\quad\text{and}$
(4) $m_{-}\bigl(t,k\bigr)\;=\;\frac{1}{k+1}\sum_{i=0}^{k}p\bigl(t-i\bigr),$
noting that $i$ runs from 0 to $k$, so there are $(k+1)$ terms in the sum. A key drawback is that the window length $k$ coincides with the prediction horizon $h$. This can bias the labels: for instance, a horizon of $h=2$ may not provide enough smoothing, whereas a large horizon might over-smooth price moves.
@@
(5) $w_{+}\bigl(t,h,k\bigr)=\frac{1}{k+1}\,\sum_{i=0}^{k}p\bigl(t+h-i\bigr)$
(6) $w_{-}\bigl(t,h,k\bigr)=\frac{1}{k+1}\,\sum_{i=0}^{k}p\bigl(t-i\bigr).$
@@
(7) $l\bigl(t,h,k\bigr)\,=\,\frac{\,w_{+}\bigl(t,h,k\bigr)\;-\;w_{-}\bigl(t,h,k\bigr)\,}{\,w_{-}\bigl(t,h,k\bigr)\!}.$
We classify a trend as upward if $l(t,h,k)>\theta$, downward if $l(t,h,k)<-\theta$, and stable if $\theta\,\leq\,l(t,h,k)\,\leq\,\theta$. The threshold $\theta$ is often chosen to balance the three classes rather than to reflect trading costs. We argue, however, that relating $\theta$ to transaction costs can better align trend predictions with profitability. Thus, in Section 6, we examine setting $\theta$ to the average spread (the difference between the best bid and ask prices) as a percentage of the mid-price [^2] Expressing the spread as a percentage of the mid-price preserves consistency with $l(t,h,k)$, which is also a percentage., since the spread represents the main transaction cost.
@@
5. Models
We propose two novel deep learning models for Stock Price Trend Prediction (SPTP) using Limit Order Book (LOB) data. The first, called MLPLOB, is a simple MLP-based model. The second, TLOB, leverages a dual-attention Transformer-based approach. Both models take as input a sequence of LOB time series consisting of the last $T$ LOB snapshots for 10 LOB levels.
5.1. MLPLOB
A key finding from the benchmark study by Prata et al. (Prata et al., 2024) reveals that, despite the proliferation of specialized deep learning architectures for SPTP, their performance often converges toward low values when tested on diverse and complex datasets. Inspired by the work of Tolstikhin et al. (Tolstikhin et al., 2021) and Zeng et al. (Zeng et al., 2023), who demonstrated that simple MLP-based models can perform as well as state-of-the-art (SoTA) methods in certain domains, we develop an MLP-based architecture for SPTP with LOB data, called MLPLOB.
Architecture Overview. MLPLOB is composed of multiple blocks, each containing two types of MLP layers:
  1. (1)
    Feature-Mixing MLPs, which operate along the feature axis.
  2. (2)
    Temporal-Mixing MLPs, which operate along the time axis.
This design aims to capture both spatial and temporal relationships in LOB data–characteristics that Sirignano and Cont (Sirignano, 2019; Sirignano and Cont, 2021) identified as fundamental to LOB dynamics and modeling.
Each MLP layer consists of two fully connected layers, mirroring the MLP component used in Transformer architectures (Vaswani et al., 2017). Initially, the input sequence is projected linearly into a tensor $\mathbf{X}\in\mathbb{R}^{T\times N}$, where $N$ is a chosen hyperparameter.

### v3

# TLOB: A Novel Transformer Model with Dual Attention for Price Trend Prediction with Limit Order Book Data

@@ Abstract
Price Trend Prediction (PTP) based on Limit Order Book (LOB) data is a fundamental challenge in financial markets. Despite advances in deep learning, existing models fail to generalize across different market conditions and assets. Surprisingly, by adapting a simple MLP-based architecture to LOB, we show that we surpass SoTA performance; thus, challenging the necessity of complex architectures. Unlike past work that shows robustness issues, we propose TLOB, a transformer-based model that uses a dual attention mechanism to capture spatial and temporal dependencies in LOB data. This allows it to adaptively focus on the market microstructure, making it particularly effective for longer-horizon predictions and volatile market conditions. We also introduce a new labeling method that improves on previous ones, removing the horizon bias. We evaluate TLOB’s effectiveness across four horizons, using the established FI-2010 benchmark, which exceeds the state-of-the-art by an average of 3.7 F1-score. Additionally, TLOB shows average improvements on Tesla and Intel with a 1.3 and 7.7 increase in F1-score, respectively. Finally, we tested TLOB on a recent Bitcoin dataset, and TLOB outperforms the SoTA performance by an average of 1.1 in F1-score. Additionally, we empirically show how stock price predictability has declined over time, -6.68 in F1-score, highlighting the growing market efficiency. Predictability must be considered in relation to transaction costs, so we experimented with defining trends using an average spread, reflecting the primary transaction cost. The resulting performance deterioration underscores the complexity of translating trend classification into profitable trading strategies. We argue that our work provides new insights into the evolving landscape of stock price trend prediction and sets a strong foundation for future advancements in financial AI. We commit to releasing the code publicly.
[…]
7943
1 Introduction
@@
Over the past few decades, the global financial landscape has undergone a profound transformation, transitioning from manual trading operations to sophisticated electronic platforms. This evolution has been so significant that by 2020, electronic trading accounted for over 99% of equity shares traded in the United States, a stark contrast to just 15% in 2000 [25]. At the heart of this revolution lies the electronic Limit Order Book (LOB), a dynamic data structure that has become the cornerstone of modern financial markets. In today’s competitive financial world, the majority of the markets utilize electronic LOBs to record trades. The continuous inflow of limit orders, organized by price levels, creates a dynamic structure that evolves over time, reflecting the real-time balance of supply and demand. However, this multidimensional structure, which spans price levels and volumes, presents complex challenges for understanding market behavior, forecasting stock price trends, and simulating realistic market conditions. The non-stationary nature of the markets, characterized by their stochastic behavior, makes modeling security price movements challenging. Traditional statistical methods fail to capture these complexities, especially when attempting to predict short-term price trends. However, recent advancements in deep learning have opened new avenues for tackling these challenges, offering the ability to model the non-linear relationships and temporal dependencies inherent in LOB data.
@@
Price Trend Prediction (PTP) [^1] In the literature, it is also referred to as mid-price movement prediction. remains one of the most challenging and economically significant problems in financial markets, attracting significant attention from academic researchers and industry practitioners. One prominent application of PTP, particularly utilizing Limit Order Book (LOB) data, lies within high-frequency trading, where algorithms attempt to capitalize on short-term price movements. Predicting future market movements is a highly challenging task due to the complexity, non-stationarity, and volatility of financial markets. However, with the growing availability of Limit Order Book (LOB) data and advancements in deep learning, new opportunities have emerged to improve the accuracy of these predictions. This paper explores the application of deep learning models to PTP using Limit Order Book (LOB) data, which provides the most granular and complete information on trades. Financial markets do not exist in a vacuum; they are continuously shaped by the actions and expectations of countless participants who, according to the Efficient Market Hypothesis (EMH), collectively incorporate all available information into asset prices. When models discover a predictive pattern and traders act on it, the anomaly is quickly competed away, causing a paradox: successful signals sow the seeds of their own demise. Over time, greater liquidity, advanced trading technologies, and the proliferation of algorithmic strategies intensify this effect, i.e., any exploitable signal becomes visible in execution data and erodes more rapidly. Consequently, the apparent decline in forecast accuracy in our findings aligns with EMH principles: as soon as new patterns are detected and exploited, the relentless engine of arbitrage drives markets back toward efficiency. This interplay underlines why forecasting often becomes more difficult the farther we move from idealized, less liquid markets (like FI-2010) toward active, high-efficiency markets (like NASDAQ), thereby illustrating a core tension between the pursuit of alpha and the self-correcting nature of competitive markets. Traditional forecasting approaches relied on technical analysis and statistical methods, but recent years have seen a shift toward more sophisticated deep learning methods. A lot of different types of deep learning architectures have been utilized to tackle the PTP tasks. Recurrent Neural Networks (RNNs) based on Long-Short Term Memory (LSTM) layers [38], Convolutional Neural Networks (CNNs) [39, 40, 45], Temporal Attention-Augmented Bilinear architecture (TABL) [36], and many others [24, 44, 30]. Recent work [31] has highlighted the limitations of existing models, particularly their lack of robustness and generalizability when applied to diverse market conditions and more efficient stocks. In this paper, we address these limitations by proposing TLOB, a transformer-based approach that outperforms all the existing models on both benchmark and real-world datasets, paving the way for more reliable PTP applications. We also introduce an MLP-based model to show that a simple architecture, based on fully connected layers and GeLU activation function, can outperform all the SoTA models. We list our contributions:
  1. 1.
@@
    Novel Architecture Proposals: We introduce two new deep learning models that surpass state-of-the-art performances:
@@
  2. 2.
    Comprehensive Evaluation: We conduct extensive experiments on the benchmark FI-2010 dataset, a NASDAQ dataset composed of Tesla and Intel stocks, and a BTC dataset, with several baselines, providing insights into model performance across different market assets, conditions, and horizons. We also perform an ablation study investigating the design choices of TLOB.
  3. 3.
@@ 2 Background
The LOB is a data structure that maintains and matches active limit orders and market orders in accordance with a predefined set of rules. This structure is transparently accessible to all market participants and is subject to continuous updates with each event, including order placement, modification, cancellation, and execution. The most widely adopted mechanism for order matching is the Continuous Double Auction (CDA) [4]. Under the CDA framework, orders are executed whenever the best bid (the highest price a buyer is willing to offer) and the best ask (the lowest price a seller is willing to accept) overlap. This mechanism facilitates continuous and competitive trading among market participants. The price of a security is commonly defined as the mid-price, calculated as the average of the best ask and best bid prices, with the difference between these prices representing the bid-ask spread.
@@
Given that limit orders are organized into distinct depth levels, each comprising bid price, bid size, ask price, and ask size, based on their respective prices, the temporal evolution of a LOB constitutes a complex, multidimensional temporal problem. Research on LOB data can be broadly categorized into four primary types: empirical analyses of LOB dynamics [12, 5], price and volatility forecasting [45, 34], stochastic modeling of LOB dynamics [13, 17], and LOB market simulation [8, 10, 26].
3 Related Work
The task of accurately modeling the complex data patterns and large volumes linked to Limit Order Books (LOBs) has driven the advancement of deep learning algorithms. In this section, we will summarize the State-of-The-Art (SoTA) deep learning models in the Price Trend Prediction (PTP) task. Tsantekidis et al. initially employed Recurrent Neural Networks (RNNs) based on Long-Short Term Memory (LSTM) layers [39] and subsequently introduced a CNN-based model (CNN) [38]. Later, they proposed CNN-LSTM [40], which combines CNN feature extraction with LSTM classification. Tran et al. developed the Temporal Attention-Augmented Bilinear Layer (TABL) [36] for multivariate time series, capturing feature and temporal dependencies via bilinear transformations. This was extended to BINCTABL [37], incorporating a bilinear normalization layer to handle non-stationarity and magnitude disparities. Passalis et al. introduced DAIN (Deep Adaptive Input Normalization) [30], a three-step adaptive normalization layer (shifting, scaling, gating), which was integrated into MLPs, CNNs, and RNNs. Zhang et al. presented DEEPLOB [45], comprising convolutional layers, an Inception Module for feature extraction, and an LSTM for temporal dependencies. They later enhanced this with an attention mechanism in DEEPLOBATT [44] for multi-horizon forecasting, using an encoder and attention-weighted hidden states. Kiesel et al. introduced Axial-LOB [24], utilizing axial attention to factorize 2D attention into separate 1D modules for feature and time axes. For a comprehensive review and benchmark of those models we refer the reader to [31].
4 Task Definition
@@
$\mathbb{L}(t)=\bigl{(}P^{ask}(t),\,V^{ask}(t),\,P^{bid}(t),\,V^{bid}(t)\bigr{)},$ (1)
where $P^{ask}(t)$ and $P^{bid}(t)\in\mathbb{R}^{L}$ are the prices at levels $1$ through $L$, and $V^{ask}(t)$ and $V^{bid}(t)\in\mathbb{R}^{L}$ are the corresponding volumes.
Trend Definition We employ a ternary classification system for price trends: U (“upward”) denotes an increasing price trend, D (“downward”) indicates a decreasing trend, and S (“stable”) represents price movements with only minor variations.
In equity markets, mid-prices are generally considered the most reliable single-value indicator of actual asset prices. However, owing to inherent market fluctuations and exogenous shocks, mid-prices can exhibit considerable volatility. Consequently, labeling consecutive mid-prices $\bigl{(}p_{t},p_{t+1}\bigr{)}$ often results in noisy labels.
Figure 1: Comparison of three labeling methods. $t$ is the current timestamp, $k$ is the smoothing window length, and $h$ is the prediction horizon. In our proposed method (c), $k$ and $h$ are defined independently, providing a more flexible and unbiased approach.
To mitigate this, many labeling strategies employ smoother mid-price functions, averaging prices over a chosen “window length” to reduce short-term noise and better reflect persistent directional moves. An example of this approach appears in [29], detailed in Section 6.1. However, as shown by Zhang et al. [45] (Fig. 2), smoothing only the future prices can lead to instability in trading signals. This instability often causes redundant trading actions and higher transaction costs. To address this, Tsantekidis et al. [38] proposed also smoothing past prices. They define:
$l(t,k)\,=\,\frac{m_{+}(t,k)\;-\;m_{-}(t,k)}{m_{-}(t,k)}\quad\text{where}$ (2)
$m_{+}\bigl{(}t,k\bigr{)}\;=\;\frac{1}{k+1}\sum_{i=0}^{k}p\bigl{(}t+i\bigr{)}\quad\text{and}$ (3)
$m_{-}\bigl{(}t,k\bigr{)}\;=\;\frac{1}{k+1}\sum_{i=0}^{k}p\bigl{(}t-i\bigr{)},$ (4)
noting that $i$ runs from 00 to $k$, so there are $(k+1)$ terms in the sum. A key drawback is that the window length $k$ coincides with the prediction horizon $h$. This can bias the labels: for instance, a horizon of $h=2$ may not provide enough smoothing, whereas a large horizon might over-smooth price moves.
@@
$w_{+}\bigl{(}t,h,k\bigr{)}=\frac{1}{k+1}\,\sum_{i=0}^{k}p\bigl{(}t+h-i\bigr{)}$ (5)
$w_{-}\bigl{(}t,h,k\bigr{)}=\frac{1}{k+1}\,\sum_{i=0}^{k}p\bigl{(}t-i\bigr{)}.$ (6)
@@
$l\bigl{(}t,h,k\bigr{)}\,=\,\frac{\,w_{+}\bigl{(}t,h,k\bigr{)}\;-\;w_{-}\bigl{(}t,h,k\bigr{)}\,}{\,w_{-}\bigl{(}t,h,k\bigr{)}\!}.$ (7)
We classify a trend as upward if $l(t,h,k)>\theta$, downward if $l(t,h,k)<-\theta$, and stable if $-\theta\,\leq\,l(t,h,k)\,\leq\,\theta$. The threshold $\theta$ is often chosen to balance the three classes rather than to reflect trading costs. We argue, however, that relating $\theta$ to transaction costs can better align trend predictions with profitability. Thus, in Section 8, we examine setting $\theta$ to the average spread (the difference between the best bid and ask prices) as a percentage of the mid-price [^2] Expressing the spread as a percentage of the mid-price preserves consistency with $l(t,h,k)$, which is also a percentage., since the spread represents the main transaction cost.
@@
Figure 2: TLOB architecture overview. The model leverages Temporal Self-Attention and Feature Self-Attention within each TLOB block to capture time-wise and spatial relationships in Limit Order Book data. Each block is preceded by Bilinear Normalization to address non-stationarity, followed by an MLPLOB block.
5 Models
We propose two novel deep learning models for Price Trend Prediction (PTP) using Limit Order Book (LOB) data. The first, called MLPLOB, is a simple MLP-based model. The second, TLOB, leverages a dual-attention Transformer-based approach. Both models take as input a sequence of LOB time series consisting of the last $T$ LOB snapshots for 10 LOB levels.
5.1 MLPLOB
A key finding from the benchmark study by Prata et al. [31] reveals that, despite the proliferation of specialized deep learning architectures for PTP, their performance often converges toward low values when tested on diverse and complex datasets. Inspired by the work of Tolstikhin et al. [35] and Zeng et al.  [43], who demonstrated that simple MLP-based models can perform as well as state-of-the-art (SoTA) methods in certain domains, we develop an MLP-based architecture for PTP with LOB data, called MLPLOB.
Architecture Overview. MLPLOB is composed of multiple blocks, each containing two types of MLP layers: (1.) Feature-Mixing MLPs, which operate along the feature axis. (2.) Temporal-Mixing MLPs, which operate along the time axis. This design aims to capture both spatial and temporal relationships in LOB data–characteristics that Sirignano and Cont [34, 33] identified as fundamental to LOB dynamics and modeling.
Each MLP layer consists of two fully connected layers, mirroring the MLP component used in Transformer architectures [41]. Initially, the input sequence is projected linearly into a tensor $\mathbf{X}\in\mathbb{R}^{T\times N}$, where $N$ is a chosen hyperparameter.